# Scaling and Predictability in Surface Quasi-Geostrophic Turbulence

### v1

@@ 4 Conclusions
Our results indicate that, for $\mathrm{Re}\gtrsim 2\times 10^{4}$, the energy spectrum approaches the Kolmogorov-like scaling $E(k)\propto k^{-5/3}$. This observation, together with the convergence in Reynolds of the prefactor $C_{K}$ suggests that SQG turbulence exhibits a well-defined inertial range—similar to that of 3D Navier-Stokes turbulence, but it does so only at very large Reynolds numbers. This regime was not observed by earlier studies at moderate Reynolds numbers.
In terms of predictability, we showed that the Lyapunov exponent scales anomalously with the Reynolds number as $\lambda\propto\text{Re}^{0.7}$, exceeding the dimensional prediction $\lambda\propto\text{Re}^{1/2}$. This anomalous scaling is reminiscent of similar behavior observed in 3D turbulence, where $\lambda\propto\text{Re}^{0.64}$, suggesting that in both systems, the deviation from dimensionality originates from a shared dynamical property. Beyond the average growth rate of infinitesimal perturbations, we investigated the statistical properties of FTLEs. Notably, both the variance and the shape of the associated Cramér function exhibit self-similar behavior across Reynolds numbers when appropriately rescaled. The ratio $\Omega/\lambda$ remains approximately constant across Re, indicating a form of universality in the core of the FTLE distribution.
Although we have studied the predictability problem from the point of view of the exponential growth of infinitesimal perturbations, it would be very interesting to investigate in detail the complementary regime of large errors and the statistics of finite-size Lyapunov exponents (Boffetta & Musacchio, 2017). Recent results have been obtained in the case of a decaying SPE cascade (Valade et al., 2024), where the authors were able to connect the hyperdiffusive behavior of Lagrangian fluid parcels with the anomalous diffusion of the system.
@@ Acknowledgement
Supported by Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations). We acknowledge HPC CINECA for computing resources within the INFN-CINECA Grants INFN24-FieldTurb and INFN25-FieldTurb.

### v2

@@ 4 Conclusions
Our results indicate that, for $\text{Re}\gtrsim 2\times 10^{4}$, the energy spectrum approaches the Kolmogorov-like scaling $E(k)\propto k^{-5/3}$. This observation, together with the convergence in Reynolds of the prefactor $C_{K}$ suggests that SQG turbulence exhibits a well-defined inertial range—similar to that of 3D Navier-Stokes turbulence, but it does so only at very large Reynolds numbers. This regime was not observed by earlier studies at moderate Reynolds numbers.
In terms of predictability, we showed that the Lyapunov exponent scales anomalously with the Reynolds number as $\lambda\propto\text{Re}^{0.7}$, exceeding the dimensional prediction $\lambda\propto\text{Re}^{1/2}$. The presence of an anomalous scaling is reminiscent of similar behavior observed in 3D Navier-Stokes turbulence, (Boffetta & Musacchio, 2017; Mohan et al., 2017; Berera & Ho, 2018; Ge et al., 2023). In particular, it has been observed that $\lambda\propto\text{Re}^{0.64}$ (Boffetta & Musacchio, 2017). This is in contrast with the scaling of the Lyapunov exponent of Lagrangian trajectories $\lambda_{L}$, which has been found to be in agreement with the Ruelle’s dimensional prediction $\lambda_{L}\propto\tau_{\nu}$ (Bec et al., 2006) up to minor corrections which can be estimated from multifractal arguments.
The difference between Eulerian and Lagrangian predictability in NS turbulence can be ascribed to the fact that the Lagrangian Lyapunov exponent is determined by the viscous time-scale $\tau_{\nu}$, while the Eulerian predictability can be influenced also by the sweeping time-scale $\tau_{E}=\ell_{\nu}/U$, where $U$ is the RMS velocity (Ge et al., 2023). Considering that $\tau_{E}\ll\tau_{\nu}$ and that $\tau_{\nu}/\tau_{E}\simeq\text{Re}^{1/4}$, the dependence of $\lambda$ on Re can be rewritten as a combination of two time scales $\lambda=\tau_{\nu}^{-1}\text{Re}^{\xi}=\tau_{\nu}^{-1+4\xi}\tau_{E}^{-4\xi}$. The Ruelle scaling is recovered for $\xi=0$, while for $\xi=1/4$ the predictability is determined by the sweeping time. Our results in SQG give $\xi^{SQG}=0.2$ which is larger than the value observed in 3D NS $\xi^{NS}=0.14$, suggesting that the sweeping effect has a stronger influence on the Eulerian predictability in SQG than in NS.
It is worth to note that the dimensional arguments discussed above do not consider the effects of intermittency corrections. It would be interesting to develop a multifractal-like approach for the direct cascade of SQG to check how the observed intermittency (Valade et al., 2025) affects the dimensional scaling of the Lyapunov exponent.
Beyond the average growth rate of infinitesimal perturbations, we investigated the statistical properties of FTLEs. Notably, both the variance and the shape of the associated Cramér function exhibit self-similar behavior across Reynolds numbers when appropriately rescaled. The ratio $\Omega/\lambda$ remains approximately constant across Re, indicating a form of universality in the core of the FTLE distribution.
Although we have studied the predictability problem from the point of view of the exponential growth of infinitesimal perturbations, it would be very interesting to investigate in detail the complementary regime of large errors and the statistics of finite-size Lyapunov exponents (Boffetta & Musacchio, 2017). Recent results have been obtained in the case of a decaying SPE cascade in SQG (Valade et al., 2024), where the authors were able to connect the hyperdiffusive behavior of Lagrangian fluid parcels with the anomalous diffusion of the system.
@@ Acknowledgement
We acknowledge HPC CINECA for computing resources within the INFN-CINECA Grants INFN24-FieldTurb and INFN25-FieldTurb.
Funding
This work was supported by Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations).
Declaration of interests
The authors report no conflict of interest.
Data availability statement
The data that support the findings of this study are available upon request.
Author ORCID
V.J. Valadão, https://orcid.org/$0$-0002-7603-5969; F. De Lillo, https://orcid.org/$0$-0002-1327-695X; S. Musacchio, https://orcid.org/$0$-0002-4564-8527; G. Boffetta,
https://orcid.org/$0$-0002-2534-7751.
Author contributions
V.J.V. performed the simulations and analyzed the data. All authors contributed equally in the discussions to reach conclusions and to write the paper.